OpenStudy (animelover91804):

What value of d makes the statement true? d + 12.6 = 9 A. –21.6 B. –3.6 C. 3.6 D. 21.6

OpenStudy (animelover91804):

XP

OpenStudy (alexandervonhumboldt2):

subtract 12.6 from both sides

OpenStudy (alexandervonhumboldt2):

d+12.6-12.6=9-12.6 d=9-12.6=???